can we find the domain of a cubic polynomial 3x-4xcube<=1




Grade:

2 Answers

Ramesh V
14 years ago

3x - 4x3 <= 1

i.e., f(x)=4x3 - 3x -1 >= 0

for x=-1, f(x)= 0 , so (x+1) is a factor , the other factors are (2x+1)

f(x)=(x+1)(2x+1)2 >=0

as (2x+1)2 is always positive

domain is (x+1)>=0

or x lies in [-1 , infininty)

--

Regards

Ramesh

vikas parashar
14 years ago

domain of any polynomial function is from -infinity to +infinity